Tohban Report 2012-02-15

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 8 Feb 2010
End Date: 15 Feb 2010
Tohban: Juan Carlos Martinez Oliveros
Tohban email:
Next Tohban: Pascal Saint-Hilaire
List all reports


Solar Activity

Solar activity was very low. There are currently 5 sunspots on disk. No substantial increase in solar activity is expected in the next week.

How many GOES flares occurred?

 Flares above B, C, M, X class were      25    15     0     0

RHESSI flare list contains 0 over the time range 08-Feb-12 -- 15-Feb-12

RHESSI Operations

Shutter and attenuator motions are been performed to study the spacecraft pointing.

It is expected that the detectors will be turn-on on February 20.


The detectors are still cooling down. Currently the cryocooler power is at 60W.

The latest numbers as of Feb. 15's at 18:55:12 UTC are:

 Cryocooler Power  		59.909	W
 Cryocooler Current  		2.227	A
 Accelerometer Average  	12.361	mg
 Cold Tip Temperature 1  	103.900	K
 Cold Tip Temperature 2  	102.025	K
 Cold Plate Temperature 1  	133.833	K
 Cold Plate Temperature 2  	131.882	K

Data Gaps

No significant data gaps during this week.


The RHESSI State-of-Health Plots are not available after 04-Feb-2012.

Personal tools